Two automobiles are equipped with the same single-frequency horn. When one is at rest and the

      other is moving toward an observer at 15ms-1, a beat frequency of 5.5Hz is heard. What is the

      frequency the horns emit? Assume T = 20°C.


1
Expert's answer
2021-03-09T15:31:11-0500

The beat frequency can be found as follows:


"f_{beat}=f'-f,"

here, "f" is the original frequency and "f'" is the Doppler shift frequency.

The Doppler shift frequency can be written as follows:


"f'=f(\dfrac{v}{v-v_s})."


Substituting "f'" into the previous equation, we get:


"f_{beat}=f(\dfrac{v}{v-v_s})-f=f[(\dfrac{v}{v-v_s})-1],""f=\dfrac{f_{beat}}{[(\dfrac{v}{v-v_s})-1]}=\dfrac{5.5\ Hz}{[(\dfrac{343\ \dfrac{m}{s}}{343\ \dfrac{m}{s}-15\ \dfrac{m}{s}})-1]}=120.3\ Hz."